# CodeRAG-Bench: Can Retrieval Augment Code Generation?

### v1

@@ Abstract
While language models (LMs) have proven remarkably adept at generating code, many programs are challenging for LMs to generate using their parametric knowledge alone. Providing external contexts such as library documentation can facilitate generating accurate and functional code. Despite the success of retrieval-augmented generation (RAG) in various text-oriented tasks, its potential for improving code generation remains under-explored. In this work, we conduct a systematic, large-scale analysis by asking: in what scenarios can retrieval benefit code generation models? and what challenges remain? We first curate a comprehensive evaluation benchmark, CodeRAG-Bench, encompassing three categories of code generation tasks, including basic programming, open-domain, and repository-level problems. We aggregate documents from five sources for models to retrieve contexts: competition solutions, online tutorials, library documentation, StackOverflow posts, and GitHub repositories. We examine top-performing models on CodeRAG-Bench by providing contexts retrieved from one or multiple sources. While notable gains are made in final code generation by retrieving high-quality contexts across various settings, our analysis reveals room for improvement—current retrievers still struggle to fetch useful contexts especially with limited lexical overlap, and generators fail to improve with limited context lengths or abilities to integrate additional contexts. We hope CodeRAG-Bench serves as an effective testbed to encourage further development of advanced code-oriented RAG methods.
@@ 1 Introduction
The task of generating program code from natural language (NL) descriptions has rapidly advanced with language models (LMs) [5, 20, 19, 34]. While more advanced code generation models are constantly emerging [23, 43, 10], most of these models employ an NL-to-code generation paradigm without the ability to integrate additional context. However, it is often challenging to directly generate programs without additional information in many complex coding scenarios, e.g., when using unfamiliar libraries that models cannot easily memorize [47, 15]. Further, solely relying on parametric knowledge learned during training also makes it harder to adapt generation to new distributions during testing [2]. For example, models are unable to stay up-to-date with continuously-evolving public libraries [47], or private code bases that are not included in the pre-training data [46, 15].
Retrieval-augmented generation (RAG) [18, 11] retrieves and incorporates relevant documents at inference time. RAG reduces the need to include all knowledge within model parameters [2], leading to accuracy improvements in various scenarios [13], even without additional training [31, 25]. It also allows for flexible knowledge updates by swapping datastores—large-scale data used during inference. Nevertheless, prior work often focuses on general-domain text-oriented generation tasks [30] using a general datastore such as Wikipedia [2]. While several works explore ways to incorporate library documents [47, 37] or files within a repository [46, 15], retrieval-augmented approaches on other types of coding problems and diverse retrieval sources are still largely underexplored.
@@
[…]
Based on this benchmark, we conduct holistic evaluations in retrieval, generation, and RACG scenarios (§3). Although code generation models can benefit from ground-truth documents in multiple scenarios, current retrieval models struggle with selecting accurate documents, especially for open-domain tasks. Meanwhile, many code generation models experience little gains due to their limited context capacity to consume retrieved documents, or limited ability to do RACG effectively.
Beyond canonical retrieval (i.e., from the ground-truth source), we also explore RACG with open retrieval, i.e., retrieving documents from various sources with different chunking strategies (§4). We find that each type of coding task can benefit from functionally relevant snippets from certain sources, and chunking documents to 200–800 tokens often gives the best results. We hope CodeRAG-Bench can serve as a testbed for future work exploring, analyzing, and improving RACG systems.
@@ 2 The CodeRAG-Bench
For CodeRAG-Bench (Figure 1), the curation methodology is motivated by the following three factors: (i) Diverse tasks: Code generation involves versatile tasks that operate on different levels (line, function, repository) and various domains (closed, open). (ii) Rigorous and reproducible evaluation: We provide high-quality annotation of ground-truth documents to enable retrieval evaluation, and execution-based evaluation for all code generation tasks to rigorously measure functional correctness. (iii) Unified interface: While current datasets utilize heterogeneous pipelines, our codebase provides a unified interface for retrieval, augmented generation, and evaluation.
In this section, we introduce the creation process of CodeRAG-Bench: programming problem integration (§2.1), retrieval source collection (§2.2), canonical document annotation (§2.3), and the evaluation pipeline (§2.4). Examples with canonical documents are available in §B.
@@ 2.1 Programming Problems
Type Dataset # Examples Ground-Truth Docs Evaluation
Basic programming HumanEval   164 program solutions execution
MBPP   500 program solutions execution
LiveCodeBench   400 - execution
Open-domain DS-1000 1000 docs execution
ODEX   945 docs, stackoverflow execution
Repository-level RepoEval (function)   373 github repository execution
SWE-bench-Lite   300 github repository execution
Code retrieval CodeSearchNet-Py 22177 CSN functions ndcg@10
Basic programming problems  This category include interview-style problems that mostly require Python built-in operations and pose algorithmic challenges. We select the two most widely used datasets: HumanEval [5] and MBPP [3], which ask the model to complete a function from an NL problem description. However, due to limited public knowledge about model training data, it is unclear whether models suffer from data contamination on HumanEval and MBPP [14]. Hence, we also include LiveCodeBench [14] with problems collected from coding websites after the training cutoff of LMs that we consider, to decrease the risk of contamination.
Open-domain problems  Open-domain coding problems require Python libraries beyond the standard libraries used in basic programming problems. We adopt the DS-1000 [17] and ODEX [42] datasets that cover data-science and general open-domain coding problems. DS-1000 collects data science problems with programs using seven common data-related libraries such as pandas and numpy. ODEX covers problems using a broader range of 79 libraries, such as web requests with requests and database operations with sqlalchemy.
Repository-level coding problems  Beyond function-level coding, some problems require editing files in the context of an entire GitHub repository. We thus adopt RepoEval [46] and SWE-bench [15] for repository-level code generation and issue-solving tasks. We integrate all three splits of RepoEval but only report its function split, as it is the only split supporting execution-based evaluation. [^2] Two other splits (API and line completion) are evaluated by lexical measures that have been shown as ineffective in signifying functional correctness [5, 42]. Notably, our codebase is the first to provide reproducible execution evaluation on RepoEval. SWE-bench focuses on resolving GitHub issues by asking models to edit multiple files that pass the required test cases. However, due to reproducibility issues on the full dataset, [^3] https://github.com/princeton-nlp/SWE-bench/issues we use SWE-bench-Lite, [^4] https://www.swebench.com/lite.html a 300-problem subset whose results can be reproduced, in addition to a packaged Docker container [^5] https://github.com/OpenDevin/OpenDevin/tree/main/evaluation/swe_bench#opendevin-swe-bench-docker-image with the pre-populated evaluation environment setup.
@@ 2.2 Retrieval Sources
Table 2: Five retrieval sources.
Resource Size Length
@@
Programming solutions  We create one document from each basic programming problems that have canonical solutions (i.e., HumanEval and MBPP), following VoyageAI [40], by concatenating its NL problem and program solution.
Online tutorials  We collect tutorials from multiple websites including GeeksforGeeks, W3Schools, tutorialspoint, and Towards Data Science, [^6] https://geeksforgeeks.org; https://www.w3schools.com/; https://www.tutorialspoint.com/; https://towardsdatascience.com via the raw HTML pages obtained from ClueWeb22 [29], a large-scale crawled web corpus. Each page contains code snippets and their text explanations, covering topics from basic programming techniques to advanced library usage.
Library documentation  We collect the official documentation provided by devdocs.io for all Python libraries following [47]. These could be especially useful for open-domain and repository-level problems that use some library functions to realize complex setups.
StackOverflow posts  StackOverflow (SO) is among the most frequently visited sites for developers. We collect all SO posts from the RedPajama-1T [7] stackexchange split. We treat each post as a retrievable document, that has a question description, code responses, and textual explanations.
GitHub repository  Lastly, we collect high-quality repositories from GitHub, using the github split of RedPajama-1T [7], as developers often refer to popular repositories when writing their own programs. Following this practical paradigm, we enable LMs to retrieve files from other GitHub repositories as contexts to write the current program.
@@ 2.3 Canonical Document Annotation
To enable reliable retrieval evaluation and estimate the upper bound of a RACG system with a perfect retriever, it is crucial that all examples include canonical documents: the document(s) containing the supporting contexts needed to solve the programming problem. However, because RACG is under-explored, most existing datasets do not provide these canonical documents. Therefore, we annotate the canonical documents from their corresponding retrieval pool, as listed in Table 1.
Basic programming problems  The canonical document for each example in HumanEval and MBPP is the document we created in §2.2 in the programming solutions pool. Since LiveCodeBench does not provide solutions to its problems, we do not annotate canonical documents for it.
@@ 2.4 Evaluation Metrics
For retrieval, we evaluate NDCG, Precision and Recall [39] and use NDCG@10 percentage as our primary metric, following prior work [12]. We only evaluate the canonical retrieval setup. For code generation, we adopt the pass@k metric [5] to measure the execution correctness of programs. We evaluate the final RAG performance both in canonical and open retrieval setups.
Table 3: Retrieval performance (NDCG@10) on code generation datasets. We do not evaluate LiveCodeBench because no ground-truth solutions are provided. RepoEval is function level with 2k context tokens. The numbers next to the method names indicate their embedding dimension size.
Method Problem Solutions Library Docs In-Repository Files
HumanEval MBPP CSN DS-1000 ODEX RepoEval SWE-bench-Lite
BM25 100.0 98.6 89.1 5.2 6.7 93.2 43.0
GIST-base ($769$) 98.0 98.0 89.9 12.0 12.1 81.2 46.8
GIST-large ($1024$) 100.0 98.9 89.6 13.6 28.0 82.9 47.8
BGE-base ($769$) 99.7 98.0 90.0 10.8 22.0 77.5 44.9
BGE-large ($1024$) 98.0 99.0 90.6 8.9 11.5 80.4 40.1
SFR-Mistral ($4096$) 100.0 99.0 - 19.3 37.1 83.8 62.7
Voyage-code ($1536$) 100.0 99.0 - 33.1 26.6 94.3 29.1
OpenAI-03 ($1536$) 100.0 98.9 - 18.2 16.5 93.0 43.3
3 Canonical RACG: Experiments and Results
We conduct baseline experiments with multiple top-performing retrieval and generation models on CodeRAG-Bench (§3.1) using canonical data sources. We report results of document retrieval (§3.2), direct NL-to-code generation (§3.3), and end-to-end RACG using retrieved context (§3.4).
@@ 3.1 Experimental Setup
Retrieval baselines  We adopt top-performing retrievers from three categories: sparse retrievers, dense retrievers with open checkpoints, and proprietary APIs. Concretely, we use BM25 [33] to represent sparse retrievers that are often robust to domain adaptations [39]. We use dense retrievers of varying sizes, namely BGE-base/large [44], GIST-base/large [35], and SFR-Embedding-Mistral [26] which are among the top of the MTEB leaderboard [28]. We use two proprietary retrieval APIs, voyage-code-2 [40] and openai-text-embedding-small-03, which are the best options with reasonable cost in our preliminary study. We also explore reranking with BGE-reranker-base [44], which reranks the top-100 openai retrieved documents before feeding into the generation models.
Generation baselines  We adopt both code-specific LMs and strong general text-oriented LMs. For code-specific LMs, we use StarCoder2 [23], CodeGemma [38], CodeLlama [34], and DeepSeekCoder [10] in various sizes. For general text LMs, we include three top-performing models: Llama3 [27], Command-R [6] specially optimized for RAG, and proprietary GPT models gpt-3.5-turbo-0125 and gpt-4. We use the instruct version of all generation models if available, since they often perform better than the base versions.
Experimental setup and hyper-parameters  For retrieval, we implement BM25 retrievers using pyserini [21] with parameter $k_{1}=1.2$ and $b=0.75$, and use sentence-transformers [32] [^7] https://sbert.net/ for all dense models with open checkpoints. For code generation, we use temperature $t=0.2$, $top\_p=0.95$ and sample one response for all generations. We prepend the top-5 retrieved documents to the original problems, and do not include other unnecessary contexts such as few-shot demonstrations. [^8] For basic programming examples that have the exact answer in the canonical documents, we remove canonical solutions from retrieved documents to avoid disclosing the answer during generation.

### v2

@@ Abstract
While language models (LMs) excel at generating code, many programs are difficult to generate using only parametric knowledge. Despite the success of retrieval-augmented generation (RAG) in text-centric tasks, its potential for code generation remains under-explored. This work introduces CodeRAG-Bench, a holistic retrieval-augmented code generation benchmark covering tasks like basic programming, open-domain, and repository-level problems and provide reproducible evaluations on both retrieval and end-to-end code generation performance. We further create a diverse, open datastore for code retrieval, aggregating sources such as competition solutions, tutorials, library documentation, StackOverflow posts, and GitHub repositories. Based on CodeRAG-Bench, we conduct large-scale evaluations of 10 retrievers and 10 LMs and systematically analyze when retrieval can benefit code generation models and identify remaining challenges. We find that while retrieving high-quality contexts improves code generation, retrievers often struggle to fetch useful contexts, and generators face limitations in using those contexts effectively. We hope CodeRAG-Bench encourages further development in code-oriented RAG methods.
[…]
CodeRAG-Bench:
Can Retrieval Augment Code Generation?
Zora Zhiruo Wangthanks: Equal contribution.  Akari Asai♢∗ Xinyan Velocity YuFrank F. Xu Yiqing Xie Graham Neubig Daniel Fried Carnegie Mellon University  University of Washington University of Southern California https://code-rag-bench.github.io/
@@ 1 Introduction
Generating code from natural language has rapidly advanced with language models (LMs; Chen et al. 2021; Li et al. 2022, 2023; Roziere et al. 2023). However, most models follow an NL (Natural Language)-to-code approach without integrating external context, which is crucial in complex scenarios like using unfamiliar libraries (Zhou et al., 2023; Jimenez et al., 2024). Relying solely on parametric knowledge also limits adaptation to new data distributions at test time, such as evolving public libraries or private code bases not seen during training Zhang et al. (2023); Jimenez et al. (2024).
Retrieval-augmented generation (RAG; Lewis et al. 2020; Guu et al. 2020) addresses this by retrieving relevant documents at inference time, reducing reliance on model parameters Asai et al. (2024) and improving accuracy across tasks Izacard et al. (2022b). Despite success in text-based tasks, its application to diverse coding problems and retrieval sources remains under-explored (Zhou et al., 2023; Su et al., 2024).
@@
We present CodeRAG-Bench, a holistic benchmark designed to advance research in retrieval-augmented code generation (RACG; §2). CodeRAG-Bench (as in Figure 1) covers six programming tasks across four categories: basic programming, open-domain coding, repository-level, and code retrieval tasks. For each task, we manually annotate canonical documents as references for evaluating RACG systems. We also compile a diverse corpus of documents from five sources: programming solutions, online tutorials, Python library documentation, StackOverflow posts, and GitHub files. In total, CodeRAG-Bench has 9k coding tasks and 25 million retrieval documents, providing a robust foundation for reproducible and reliable evaluations in retrieval and RACG.
We conduct holistic evaluations in retrieval, generation, and RACG (§3). Code generation models significantly benefit from access to canonical documents (i.e., from the canonical retrieval corpus) in various scenarios. For example, GPT-4o achieves a 27.4% gain on SWE-Bench and a 6.9% gain on the harder ODEX subset when canonical documents are provided. In RACG settings, where models retrieve top relevant documents, some even surpass their performance when using gold documents, highlighting the strong potential of retrieval-augmented approaches for enhancing code generation. However, current retrieval models face challenges in selecting useful documents, particularly for open-domain and repository-level tasks. Additionally, generation models with limited context windows exhibit smaller improvements, suggesting considerable room for future advancements.
Beyond canonical retrieval, we also explore RACG with open retrieval, i.e., retrieving documents from various sources with different chunking strategies (§4). We find that models can benefit from functionally relevant snippets from certain sources, and chunking documents to 200–800 tokens often gives the best results. For instance, by retrieving from StackOverflow or online tutorials, both StarCoder and GPT4o can significantly improve, while on repository-level tasks, the gains are rather limited. Overall, we hope CodeRAG-Bench can serve as a testbed for future work exploring, analyzing, and improving RACG systems.
@@ 2 The CodeRAG-Bench
For CodeRAG-Bench (Figure 1), the curation is driven by three factors: (i) Diverse tasks: Code generation spans multiple levels (line, function, repository) across closed and open domains. (ii) Rigorous evaluation: We offer high-quality ground-truth annotations for retrieval and execution-based evaluation to measure functional correctness. (iii) Unified interface: Our codebase provides a consistent interface for retrieval, augmented generation, and evaluation, unlike current datasets with varied pipelines.
In this section, we introduce the creation process of CodeRAG-Bench: programming problem integration (§2.1), retrieval source collection (§2.2), canonical document annotation (§2.3), and the evaluation pipeline (§2.4). Examples with canonical documents are available in §A.
Type Dataset # Examples # Corpus Ground-Truth Docs Evaluation
Basic programming HumanEval   164   164 program solutions execution
MBPP   500   500 program solutions execution
LiveCodeBench   400 - - execution
Open-domain DS-1000 1000 34,003 docs execution
ODEX   945 34,003 docs, stackoverflow execution
Repository-level RepoEval (function)   373   237 github repository execution
SWE-bench-Lite   300 40,868 github repository execution
Code retrieval CodeSearchNet-Py 22,177 22177 CSN functions ndcg@10
@@ 2.1 Programming Problems
Basic programming problems  This category includes interview-style problems that mostly require Python built-in operations and pose algorithmic challenges. We select the two most widely used datasets: HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), which ask the model to complete a function from an NL problem description. However, due to limited public knowledge about model training data, it is unclear whether models suffer from data contamination on HumanEval and MBPP (Jain et al., 2024). Hence, we also include LiveCodeBench (Jain et al., 2024) with problems collected from coding websites after the training cutoff of LMs that we consider, to decrease the risk of contamination.
Open-domain problems  Open-domain coding problems require Python libraries beyond the standard libraries used in basic programming problems. We adopt the DS-1000 (Lai et al., 2023) and ODEX (Wang et al., 2023b) datasets that cover data-science and general open-domain coding problems. DS-1000 collects data science problems with programs using seven common data-related libraries such as pandas and numpy. ODEX covers problems using a broader range of 79 libraries, such as web requests with requests and database operations with sqlalchemy.
Repository-level coding problems  Beyond function-level, some problems require editing files in the context of an entire GitHub repository. We thus adopt RepoEval (Zhang et al., 2023) and SWE-bench (Jimenez et al., 2024) for repository-level code generation and issue-solving tasks. We integrate all three splits of RepoEval but only report its function split, as it is the only split supporting execution-based evaluation. [^2] Two other splits (API and line) are evaluated by lexical measures that have been shown as ineffective in signifying functional correctness (Chen et al., 2021; Wang et al., 2023b). Notably, our codebase is the first to enable reproducible execution evaluation on RepoEval. SWE-bench focuses on resolving GitHub issues by asking models to edit multiple files that pass the required test cases. We use SWE-bench-Lite, [^3] https://www.swebench.com/lite.html a 300-problem subset whose results can be reproduced, with a packaged Docker container (Wang et al., 2024).
@@ 2.2 Retrieval Sources
Programming solutions  We create one document from each basic programming problems that have canonical solutions (i.e., HumanEval and MBPP), following VoyageAI (2024), by concatenating its NL problem and program solution.
Online tutorials  We collect tutorials from multiple websites including GeeksforGeeks, W3Schools, tutorialspoint, and Towards Data Science, [^4] https://geeksforgeeks.org; https://www.w3schools.com/; https://www.tutorialspoint.com/; https://towardsdatascience.com via the raw HTML pages obtained from ClueWeb22 Overwijk et al. (2022), a large-scale crawled web corpus. Each page contains code snippets and their text explanations, covering topics from basic programming techniques to advanced library usage.
Library documentation  We collect the official documentation provided by devdocs.io for all Python libraries following (Zhou et al., 2023). These could be especially useful for open-domain and repository-level problems that use some library functions to realize complex setups.
StackOverflow posts  StackOverflow (SO) is among the most frequently visited sites for developers. We collect all SO posts from the RedPajama-1T Computer (2023) stackexchange split. We treat each post as a retrievable document, that has a question, code responses, and textual explanations.
GitHub repository  We collect high-quality repositories from GitHub, using the github split of RedPajama-1T Computer (2023), as developers often refer to popular repositories when writing their programs. Following this practical paradigm, we enable LMs to retrieve files from other repositories as contexts to write the current program.
Resource Corpus size Avg. length
@@
Table 2: Five sources to form our retrieval datastore.
@@ 2.3 Canonical Document Annotation
To ensure reliable retrieval evaluation and estimate the upper bound of a RACG system with an ideal retriever, it’s essential that all examples include canonical documents–the documents containing the necessary context to solve the programming problem. As most existing datasets lack these canonical documents, we annotate them from the corresponding retrieval pool, as shown in Table 1.
Basic programming problems  The canonical document for examples in HumanEval and MBPP is the documents we created in §2.2 in the programming solutions pool. Since LiveCodeBench does not provide solutions to its problems, we do not annotate canonical documents for it.
@@ 2.4 Evaluation Metrics
For retrieval, we evaluate NDCG, Precision and Recall Thakur et al. (2021) and use NDCG@10 percentage as our primary metric, following prior work Izacard et al. (2022a). For code generation, we adopt the pass@k metric (Chen et al., 2021) to measure the execution correctness of programs. We evaluate the final RAG performance both in canonical and open retrieval setups.
Method Problem Solutions Library Docs In-Repository Files Avg.
HumanEval MBPP CSN DS-1000 ODEX RepoEval SWE-bench-Lite All
BM25 100.0 98.6 89.1 5.2 6.7 93.2 43.0 57.7
GIST-base ($768$) 98.0 98.0 89.9 12.0 12.1 81.2 46.8 58.0
GIST-large ($1024$) 100.0 98.9 89.6 13.6 28.0 82.9 47.8 61.7
BGE-base ($768$) 99.7 98.0 90.0 10.8 22.0 77.5 44.9 58.8
BGE-large ($1024$) 98.0 99.0 90.6 8.9 11.5 80.4 40.1 56.3
SFR-Mistral ($4096$) 100.0 99.0 - 19.3 37.1 83.8 62.7 67.0
Codesage-small ($768$) 100.0 96.3 90.7 8.9 14.3 94.1 47.1 60.1
Jina-v2-code ($768$) 100.0 97.7 - 26.2 19.9 90.5 58.3 65.4
OpenAI-03 ($1536$) 100.0 98.9 - 18.2 16.5 93.0 43.3 61.7
Voyage-code ($1536$) 100.0 99.0 - 33.1 26.6 94.3 29.1 63.7
Table 3: Retrieval performance (NDCG@10) on code generation datasets. LiveCodeBench is excluded due to lack of ground-truth solutions. RepoEval is at the function level with 2k context tokens. Embedding dimension sizes are listed next to method names. Bold indicates best performance, underline indicates second-best. Highlighted models are specifically trained for code domains. Avg. reflects overall scores, excluding CodeSearchNet.
3 Canonical RACG
We evaluate 10 top retrieval and 10 generation models on CodeRAG-Bench with canonical data sources. We report results of document retrieval (§3.2), direct NL-to-code generation (§3.3), and end-to-end RACG with retrieved context (§3.4).
@@ 3.1 Experimental Setup
Retrieval baselines  We adopt 10 top-performing retrievers from three categories: sparse, dense, and proprietary APIs. For sparse retrievers, we use BM25 Robertson and Zaragoza (2009), known for its robustness in domain adaptation Thakur et al. (2021). Dense retrievers include BGE-base/large (Xiao et al., 2023), GIST-base/large (Solatorio, 2024), and SFR-Embedding-Mistral (Meng et al., 2024), all top-ranked on the MTEB leaderboard (Muennighoff et al., 2022). We also include open code embedding models, Codesage-small Zhang et al. (2024) and Jina-v2-code Günther et al., 2023, which are specifically trained for code retrieval. Proprietary APIs include voyage-code-2 (VoyageAI, 2024), optimized for code retrieval, and openai-text-embedding-small-03, selected for its cost-effectiveness. Finally, we apply reranking with BGE-reranker-base(Xiao et al., 2023) on top-100 openai results before generation.
Generation baselines  We adopt both code-specific LMs and strong general text-oriented LMs. For code-specific LMs, we use StarCoder2 (Lozhkov et al., 2024), CodeGemma (Team, 2024), CodeLlama (Roziere et al., 2023), and DeepSeekCoder (Guo et al., 2024) in various sizes. For general text LMs, we include three top-performing models: Llama3 (Meta, 2024), Command-R (CohereAI, 2024) specially optimized for RAG, and proprietary GPT models gpt-3.5-turbo-0125 and gpt-4o. We use the instruct version of all generation models if available, since they often perform better than the base versions.
Experimental setup  For retrieval, we implement BM25 retrievers using pyserini (Lin et al., 2021) with parameter $k_{1}=1.2$ and $b=0.75$, and use sentence-transformers Reimers and Gurevych (2019) [^5] https://sbert.net/ for all dense models with open checkpoints. We prepend the top-5 retrieved documents to the original problems (we study the number of documents in §E), and do not include other contexts such as few-shot examples. For code generation, we use temperature $t=0.2$, $top\_p=0.95$ and sample one response for all generations, following prior work (Li et al., 2023). Specifically on SWE-bench-Lite, we adopt the $n=21$ way sampling and majority-vote reranking strategy proposed by Agentless (Xia et al., 2024). [^6] We found that without these approaches, performance even with state-of-the-art GPT4o remains around 1-2%.